### v1

@@ 1 Introduction
Titanium is the most commonly used base metal for manufacturing PTLs and BPPs due to its excellent corrosion resistance. However, it is expensive, difficult to machine into complex flow-field designs, and susceptible to hydrogen embrittlement, which can cause mechanical failure. Over time, titanium also develops a thick oxide layer, which increases interfacial contact resistance (ICR) at cell interfaces and thereby reduces electrolysis efficiency. To address these challenges, precious metal coatings such as Pt or Au are often applied. This significantly increases the cost of already expensive Ti PTLs and BPPs, [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] making up roughly 50–70% of the total stack costs [10, 17].
Stainless steel presents a much more affordable and easily machinable alternative to titanium, and it offers substantially higher resistance to hydrogen embrittlement [7, 12, 13, 14, 15, 17, 18]. However, in addition to oxide layer formation, dissolved metal cations, primarily $\mathrm{Fe}^{3+}$, can leach into the system and poison the catalyst and electrolyte membrane, leading to increased electrode overpotential and reduced stack lifespan [17, 19, 20, 21].
@@
In this study, we directly address this challenge by developing 200 nm-thick, highly corrosion-resistant, conductive and compact bilayer Ti–Nb–O coatings deposited by HiPIMS on stainless steel sheets, systematically tailoring their composition over a broad range. The optimized coatings enabled DOE-compliant performance while reducing Pt loading by up to two orders of magnitude compared to conventional approaches. This combination of ultra-low Pt loading and high corrosion resistance underscores the potential of these coatings as a durable, cost-effective solution for PEM electrolyzers.
@@ 2.1 Coating preparation
Each coated stainless steel sheet was cut in half, with one half subsequently coated with a 5 nm metallic Pt overlayer via RF sputtering. A Pt target with 99.99% purity was used for the deposition, and a bare Si substrate was included in each run to verify the Pt overlayer thickness. The depositions were performed in a Leybold-Heraeus LH Z400 multitarget system, featuring a stainless steel chamber evacuated using a turbomolecular pump supported by a scroll pump. The target-to-substrate distance was maintained at 70 mm, with a base pressure below 5 mPa. The target was operated at 15 W, achieving the desired 5 nm thickness after 15 minutes of sputter deposition.
2.2 Elemental composition and structure
@@ 3.2 Electrochemical and ICR measurements
Figure 7: Potentiodynamic polarization corrosion current density ($J_{\mathrm{corr}}$) and interfacial contact resistance (ICR) of coatings deposited on stainless steel substrates under various conditions. The indicated Nb content in the metal fraction, $\mathrm{Nb}/(\mathrm{Ti}+\mathrm{Nb})$, and O content apply only to the upper layer, which is in direct contact with the electrolyte. Panels (a) and (c) show $J_{\mathrm{corr}}$ and ICR measured before the potentiostatic polarization, while panels (b) and (d) present the values after the potentiostatic polarization. The values are visualized using continuous color mapping across the compositional space of the upper layer. Data smoothing is applied to highlight overall trends within each panel. Triangle, diamond, circle, and square symbols indicate the elemental compositions of coatings with upper layers deposited at $p_{\mathrm{ox}}$ values of 0, 3, 5, and 8 mPa, respectively. The complete set of data is provided in Supplementary Table 1a. The complete set of data is provided in Supplementary Table 1a.
For the ICR before the potentiostatic polarization, most coatings with upper layers deposited at $p_{\mathrm{ox}}$ of 0 and 3 mPa exhibited ICR values below 100 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$. Notably, coatings with intermediate Nb content and an upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ demonstrated the lowest initial ICR, falling below 7 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$. These results correlate with the higher $J_{\mathrm{corr}}$ values before the potentiostatic polarization for these coatings, indicating higher surface conductivity. This can be attributed to a substoichiometric cubic TiOx phase with higher conductivity in the upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (see Fig. 4), which is, however, capable of limiting surface passivation upon exposure to air. In contrast, coatings with upper layers deposited at $p_{\mathrm{ox}}$ of 5 and 8 mPa generally showed ICR values above 200 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ (Fig. 7c). After the potentiostatic polarization, ICR values for all coatings increased substantially, exceeding $500~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ (Fig. 7d), due to increased surface oxidation that impeded effective contact with the GDL. This is consistent with the observed increase in $E_{\mathrm{corr}}$ for all coatings, which converged to a narrow range. Notably, ICR values tended to increase with both higher O and Nb content in the upper layers, reaching a maximum of 900 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, which again correlates with the $J_{\mathrm{corr}}$ trend observed after the potentiostatic polarization.
@@ 3.3 Electrochemical and ICR measurements after Pt overlayer application
The ICR values of all coatings measured before the potentiostatic polarization dropped below $50~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ after being coated with a 5 nm Pt overlayer (Fig. 9c). This large drop is not caused by decreased resistivity of the coatings, but by the increased effective contact area between the GDL wires and the upper layer provided by the thin, but highly conductive Pt layer. The lowest pre-corrosion ICR values were observed for Ti–O and Ti–Nb–O coatings deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$, with values below $5~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, except for the coating with the highest Nb content. Unexpectedly, Ti–Nb–O coatings with the highest and second-highest Nb contents deposited at $p_{\mathrm{ox}}=8~\mathrm{mPa}$ also exhibited ICR values below $5~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, deviating from the general trend.
@@
Following the potentiostatic polarization, ICR values remained largely stable, with only minor variations (see Fig. 9d). For several compositions, a slight decrease in ICR was observed after polarization, particularly for Ti–Nb–O coating deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$ with the lowest Nb content (from 4.9 to $4~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$), for the Ti–O coating with upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (from 7.6 to $5.6~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$), and for Ti–Nb–O coating deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$ with the lowest Nb content and upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (from 7.3 to $5.3~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$). Notably, these coatings also exhibited some of the lowest ICR values after the potentiostatic polarization without the Pt overlayer (see Fig. 7d). For all remaining coatings, the ICR either remained similar or increased after the potentiostatic polarization. It should be noted that deposition of the Pt overlayer within a single uninterrupted process—i.e., without exposing the upper layer to air—would likely result in even lower ICR values due to minimization of interfacial oxidation or contamination.
@@
Overall, these findings demonstrate that the developed coatings enable the DOE ICR target to be met after accelerated corrosion test with a Pt overlayer as thin as 5 nm, resulting in a significant reduction in Pt loading.
@@ 4 Conclusions
This substantial reduction in Pt loading, combined with the high corrosion resistance of the Ti–Nb–O coatings, highlights their potential as a durable, cost-effective, and sustainable solution for PEM electrolyzers with significantly reduced reliance on precious metals.

### v2

@@ 1 Introduction
Titanium is the most commonly used base metal for manufacturing PTLs and BPPs due to its excellent corrosion resistance. However, it is expensive, difficult to machine into complex flow-field designs, and susceptible to hydrogen embrittlement, which can cause mechanical failure. Over time, titanium also develops a thick oxide layer, which increases interfacial contact resistance (ICR) at cell interfaces and thereby reduces electrolyzer efficiency. To address these challenges, precious metal coatings such as Pt or Au are often applied. This significantly increases the cost of already expensive Ti PTLs and BPPs, [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] making up roughly 50–70% of the total stack costs [10, 17].
Stainless steel presents a much more affordable and easily machinable alternative to titanium, and it offers substantially higher resistance to hydrogen embrittlement [7, 12, 13, 14, 15, 17, 18]. However, in addition to oxide layer formation, dissolved metal cations, primarily $\mathrm{Fe}^{2+}$, can leach into the system and poison the catalyst and electrolyte membrane, leading to increased electrode overpotential and reduced stack lifespan [17, 19, 20, 21].
@@
In this study, we directly address this challenge by developing 200 nm-thick, highly corrosion-resistant, conductive and compact bilayer Ti–Nb–O coatings deposited by HiPIMS on stainless steel sheets, systematically tailoring their composition over a broad range. The optimized coatings enabled DOE-compliant performance while reducing Pt loading by up to two orders of magnitude compared to conventional approaches.
Although all coatings were deposited and evaluated on flat stainless steel sheets, the process and properties were developed to enable potential application to both BPPs and PTLs. For Ti PTLs, our coating protocol (including the initial etching step) could be directly implemented to modify contact surfaces on both the BPP and MEA sides. For steel PTLs, however, additional studies are needed to ensure that harmful ions do not leach from uncoated internal surfaces and compromise membrane durability. In such cases, a preliminary treatment to fully cover all internal surfaces, such as Ti electroplating, may be required before applying our coating. By outlining these considerations, we aim to motivate further research to confirm performance on actual PTL architectures and to encourage in situ testing and validation of the presented coating system in operating electrolyzers.
@@ 2.1 Coating preparation
Each coated stainless steel sheet was cut in half, with one half subsequently coated with a 5 nm metallic Pt overlayer via RF sputtering. All depositions were performed in a Leybold-Heraeus LH Z400 multitarget system, featuring a stainless steel chamber evacuated by a turbomolecular pump supported by a scroll pump. A Pt target with 99.99% purity was used, with the target-to-substrate distance maintained at 70 mm and a base pressure below 5 mPa. The target was operated at 15 W, achieving the desired 5 nm thickness after 15 minutes of sputter deposition. To verify the Pt overlayer thickness, a bare Si substrate was included in each run. For this measurement, a small drop of nail polish was applied to the Si substrate prior to deposition and removed afterward, creating a sharp step between coated and uncoated regions. The step height, corresponding to the Pt thickness, was measured using atomic force microscopy (AFM).
2.2 Elemental composition, structure, and electrical resistivity
@@
The electrical resistivity of upper layers, approximately 500 nm thick and deposited on glass substrates, was measured using a four-point probe system (Jandel MACOR Probe Head) in a linear configuration.
@@ 3.2 Electrochemical and ICR measurements
Figure 7: Potentiodynamic polarization corrosion current density ($J_{\mathrm{corr}}$) and interfacial contact resistance (ICR) of coatings deposited on stainless steel substrates under various conditions. The indicated Nb content in the metal fraction, $\mathrm{Nb}/(\mathrm{Ti}+\mathrm{Nb})$, and O content apply only to the upper layer, which is in direct contact with the electrolyte. Panels (a) and (c) show $J_{\mathrm{corr}}$ and ICR measured before the potentiostatic polarization, while panels (b) and (d) present the values after the potentiostatic polarization. The values are visualized using continuous color mapping across the compositional space of the upper layer. Data smoothing is applied to highlight overall trends within each panel. Triangle, diamond, circle, and square symbols indicate the elemental compositions of coatings with upper layers deposited at $p_{\mathrm{ox}}$ values of 0, 3, 5, and 8 mPa, respectively. The complete set of data is provided in Supplementary Table 1a.
The complete set of data is provided in Supplementary Table 1a.
@@
For the ICR before the potentiostatic polarization, most coatings with upper layers deposited at $p_{\mathrm{ox}}$ of 0 and 3 mPa exhibited ICR values below 100 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$. Notably, coatings with intermediate Nb content and an upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ demonstrated the lowest initial ICR, falling below 7 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$. These results correlate with the higher $J_{\mathrm{corr}}$ values before the potentiostatic polarization for these coatings, indicating higher surface conductivity. This can be attributed to a substoichiometric cubic TiOx phase with high conductivity in the upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (see Fig. 4). Unlike the metallic $\alpha$-Ti phase, which readily forms a surface-passivating non-conductive TiO2 oxide layer upon exposure to air, the substoichiometric TiOx phase tends to incorporate oxygen into its bulk to increase its stoichiometry, suppressing resistive surface TiO2 layer formation and maintaining lower ICR. In contrast, coatings with upper layers deposited at $p_{\mathrm{ox}}$ of 5 and 8 mPa generally showed ICR values above 200 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ (Fig. 7c), as the increased oxygen content in the substoichiometric TiOx phase in the upper layer leads to its higher resistivity (see Fig. 5), likely impeding efficient electron transfer at the contact with the GDL wires.
After the potentiostatic polarization, ICR values for all coatings increased substantially, exceeding 500 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ (Fig. 7d), due to increased surface oxidation that impeded effective contact with the GDL. This is consistent with the observed increase in $E_{\mathrm{corr}}$ for all coatings, which converged to a narrow range. Notably, ICR values tended to increase with both higher O and Nb content in the upper layers, reaching a maximum of 900 $\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, which again correlates with the $J_{\mathrm{corr}}$ trend observed after the potentiostatic polarization.
@@ 3.3 Electrochemical and ICR measurements after Pt overlayer application
The ICR values of all coatings measured before the potentiostatic polarization dropped below $50~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$ after being coated with a 5 nm Pt overlayer (Fig. 9c). This substantial decrease is not due to a reduction in the bulk resistivity of the coatings, but is primarily attributed to enhanced electron transfer between the GDL wires and the upper layer. The Pt overlayer provides a metallic interface with the GDL wires and increases the effective contact area with the upper layer. This effect is particularly significant for upper layers with higher initial electrical resistivity, where electron transfer at the contact is otherwise limited.
@@
The lowest pre-corrosion ICR values were observed for Ti–O and Ti–Nb–O coatings deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$, with values below $5~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, except for the coating with the highest Nb content. Unexpectedly, Ti–Nb–O coatings with the highest and second-highest Nb contents deposited at $p_{\mathrm{ox}}=8~\mathrm{mPa}$ also exhibited ICR values below $5~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$, deviating from the general trend. Some deviations in ICR values after Pt deposition, measured prior to potentiostatic polarization, may arise from incomplete or non-uniform Pt coverage, especially on rough or porous surfaces. Notably, industry-relevant deposition of the Pt overlayer in a single uninterrupted process (without exposing the upper layer to air) would likely yield even lower and more consistent ICR values prior to polarization, as this approach minimizes interfacial oxidation and contamination.
The most relevant results for the long-term stability and durability of the coatings under operational conditions are those obtained after the potentiostatic polarization test. As shown in Fig. 9d, ICR values remained largely stable, with only minor variations. For several compositions, a slight decrease in ICR was observed after polarization, particularly for Ti–Nb–O coating deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$ with the lowest Nb content (from 4.9 to $4~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$), for the Ti–O coating with upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (from 7.6 to $5.6~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$), and for Ti–Nb–O coating deposited at $p_{\mathrm{ox}}=0~\mathrm{mPa}$ with the lowest Nb content and upper layer deposited at $p_{\mathrm{ox}}=3~\mathrm{mPa}$ (from 7.3 to $5.3~\mathrm{m}\Omega\text{\textperiodcentered}\mathrm{cm}^{2}$). Notably, these coatings also exhibited some of the lowest ICR values after the potentiostatic polarization without the Pt overlayer (see Fig. 7d). For all remaining coatings, the ICR either remained similar or increased after the potentiostatic polarization.
Although the 5 nm Pt overlayer provides substantial initial ICR improvement, it only partially limits oxygen diffusion and does not fully prevent oxygen penetration into the underlying coating. Therefore, the intrinsic properties of the bare coating without the Pt overlayer—determined by its composition and structure, as discussed in Sec. 3.2—remain critical for long-term resistance to passivation and degradation, ensuring durable low ICR and sustained performance during operation.
@@
Overall, these findings demonstrate that the developed coatings enable the DOE ICR target to be met after accelerated corrosion testing with a Pt overlayer as thin as 5 nm, resulting in a significant reduction in Pt loading.
@@ 4 Conclusions
Our findings highlight the potential of these coatings as a durable, cost-effective, and sustainable solution for PEM electrolyzers with significantly reduced reliance on precious metals. While the present work demonstrates performance on flat stainless steel sheets, the coating strategy was developed with broader applicability in mind, motivating future research on adapting and validating this approach for more complex BPP and PTL geometries in operating electrolyzers. For steel PTLs, additional studies should focus on testing whether preliminary treatment to fully cover all internal surfaces with a protective layer is required.